OpenStudy (anonymous):

A car of mass 1500 kg has a constant speed of 40 m/s when travelling up a hill inclined at -1 1 sin 15 to the horizontal with the engine’s power is 50 kW. The car moves down the same hill at the speed of 25 m/s while the engine working at 30 kW. Find the acceleration of the car when it moves downhill if given resistance, v2 R= k where k is a constant and v is the velocity of car.help me !this is urgent .

OpenStudy (anonymous):

I am not clear about the inclination part, there is an error in the typing.

OpenStudy (anonymous):

is it ?

OpenStudy (anonymous):

A car of mass 1500 kg has a constant speed of 40 m/s when travelling up a hill inclined at -1 1 sin 15. what is -1 1 sin15 its gibberish

OpenStudy (anonymous):

\[\sin^{-1} \]1/15

OpenStudy (anonymous):

v2 is v^2

OpenStudy (anonymous):

what is v^2 R= k

OpenStudy (anonymous):

R=(v^2)/k k is constant

OpenStudy (anonymous):

50kW means 50000J/s Means in one second, it generates 50000 kinetic energy, 50 000= 0.5*m*v^2 v= 8.45m/s meaning in one second it would accelerate by 8.45m/s but it doesnt increase speed because of 2 things, one is the resistance and 2nd is gravity. Force generated by engine = 1400*8.45=11830 N Force generated by engine= resistance (we want to find this) + gravity. 11830 - (1400)(9.81 sin 6degrees)=10394 N R= 10394=v^2/k 10394= (40^2)/k k=0.154 Now moving downhill. Resistance = (25^2)/ 0.154=4058N Force due to gravity acting downhill now = 1400*9.81*sin6degrees=1435.6 N Force due to engine, 30 000 = 0.5*1400* v^2 acceleration= 6.55m/s^2 Force = 1400*6.55=9165 N Now putting all the force together, Resultant force = Force by engine+ Force by gravity - Force by resistance. = 9165+1435.6-4058 =6542.6N Acceleration= 6542.6/1400=4.67 m/s^2 Okay this is very long and doing it on the computer may make me careless, the method is right, just double check the calculation. Im tired after doing so many problems, so do forgive if there are any errors

OpenStudy (anonymous):

oh thanx ! (':